63c7= c=63!/(63-7)!63!
Combination Formula nCr=n!/(n-r)!r!

Answer by Tatiana_Stebko(1539)

nCr=n!/((n-r)!r!)
63C7= 63!/((63-7)!7!)=63!/(56!7!)=(56!*57*58*59*60*61*62*63)/(56!*1*2*3*4*5*6*7)=(57*58*59*60*61*62*63)/(2*3*4*5*6*7)=553270671